Two different isosceles triangles with perimeter 4a+b?

Is this a right angle isosceles triangle?
Answer #1
If there are two triangles, why do refer to one of them as "this"?

If a and b are two sides and the perimeter is 4a + b, the third side is
c = (4a + b) - a - b = 3a. Since c is not equal to a and the triangle is isosceles, either b = c or b = a.

Use the Pythagorean theorem to see if the triangle could be a right angle.
if a = b and c = 3a, it can't be, since a^2 + b^2 = 2a^2 cannot = c^2.
answered by: Payton